\begin{thebibliography}{KMV02}

\bibitem[HR92]{HejhalRackner1992}
D.~Hejhal and B.~Rackner, \emph{On the topography of {M}aass waveforms for
  {$\mathrm{PSL}(2, \mathbf{Z})$}}, Experiment. Math. \textbf{1} (1992), no.~4,
  275--305.

\bibitem[HS10]{HolowinskySoundararajan2010}
R.~Holowinsky and K.~Soundararajan, \emph{Mass equidistribution for {H}ecke
  eigenforms}, Ann. of Math. (2) \textbf{172} (2010), no.~2, 1517--1528.

\bibitem[KK19]{KanekoKoyama2019}
I.~Kaneko and S.~Koyama, \emph{Quantum unique ergodicity for {E}isenstein
  series in the level aspect over function fields}, in preparation, 2019.

\bibitem[KMV02]{KowalskiMichelVanderKam2002}
E.~Kowalski, P.~Michel, and J.~VanderKam, \emph{{R}ankin-{S}elberg
  {$L$}-functions in the level aspect}, Duke Math. J. \textbf{114} (2002),
  no.~1, 123--191.

\bibitem[Koy09]{Koyama2009}
S.~Koyama, \emph{Equidistribution of {E}isenstein series in the level aspect},
  Commun. Math. Phys. \textbf{289} (2009), no.~3, 1131--1150.

\bibitem[LS95]{LuoSarnak1995}
W.~Luo and P.~Sarnak, \emph{Quantum ergodicity of eigenfunctions on
  {$\mathrm{PSL}_{2}(\mathbf{Z}) \backslash \mathbf{H}^{2}$}}, Publ. Math.
  Inst. Hautes {\'{E}}tudes Sci. \textbf{581} (1995), no.~1, 207--237.

\bibitem[MV10]{MichelVenkatesh2010}
P.~Michel and A.~Venkatesh, \emph{The subconvexity problem for
  {$\mathrm{GL}_{2}$}}, Publ. Math. Inst. Hautes {\'{E}}tudes Sci. \textbf{111}
  (2010), no.~1, 171--271.

\bibitem[Nel11]{Nelson2011}
P.~D. Nelson, \emph{Equidistribution of cusp forms in the level aspect}, Duke
  Math. J. \textbf{160} (2011), no.~3, 467--501.

\bibitem[NPS14]{NelsonPitaleSaha2014}
P.~D. Nelson, A.~Pitale, and A.~Saha, \emph{Bounds for {R}ankin-{S}elberg
  integrals and quantum unique ergodicity}, J. Amer. Math. Soc. \textbf{27}
  (2014), no.~1, 147--191.

\bibitem[RS94]{RudnickSarnak1994}
Z.~Rudnick and P.~Sarnak, \emph{The behaviour of eigenstates of arithmetic
  hyperbolic manifolds}, Commun. Math. Phys. \textbf{161} (1994), no.~1,
  195--213.

\bibitem[Spi03]{Spinu2003}
F.~Spinu, \emph{The {$L^{4}$}-norm of the {E}isenstein series}, Ph.D. thesis,
  Princeton University, Princeton, New York, 2003, ProQuest LLC, Ann Arbor, MI,
  2003, \url{http://www.math.jhu.edu/~fspinu/math/thesis.pdf}.

\bibitem[You16]{Young2016}
M.~P. Young, \emph{The quantum unique ergodicity conjecture for thin sets},
  Adv. Math. \textbf{286} (2016), no.~2, 958--1016.

\bibitem[Zha19]{Zhang2019}
L.~Zhang, \emph{Quantum unique ergodicity of degenerate {E}isenstein series on
  {$\mathrm{GL}(n)$}}, Commun. Math. Phys. \textbf{369} (2019), no.~1, 1--48.

\end{thebibliography}

\documentclass[reqno,dvipdfmx]{amsart}

\newtheorem*{lemma24}{Lemma 2.4}
\newtheorem*{theorem13}{Theorem 1.3}

\theoremstyle{definition}

\numberwithin{equation}{section}

\usepackage{amsmath,amssymb,amsfonts,amsthm}
\usepackage{enumerate}
\usepackage{url}
\usepackage[breaklinks=true,colorlinks=true,linkcolor=blue,citecolor=blue,filecolor=blue,urlcolor=blue]{hyperref}
\hypersetup{linktocpage}
\usepackage{bm}
\usepackage[hiresbb]{graphicx,xcolor}
\usepackage{tcolorbox}
\usepackage{booktabs,tabularx}
\usepackage{mathtools}
\mathtoolsset{showonlyrefs=true}
\usepackage{mathrsfs}
\usepackage{float}
\usepackage{subcaption}
\usepackage{yfonts}
\usepackage{accents}
\usepackage{here}
\usepackage[top=30truemm,bottom=30truemm,left=25truemm,right=25truemm]{geometry}

\makeatletter
\newcommand{\tpmod}[1]{{\@displayfalse\pmod{#1}}}
\makeatother
\def\rnum#1{\expandafter{\romannumeral #1}} 
\def\Rnum#1{\uppercase\expandafter{\romannumeral #1}}
\makeatletter
\let\amsmath@bigm\bigm
\renewcommand{\bigm}[1]{%
\ifcsname fenced@\string#1\endcsname
\expandafter\@firstoftwo
\else
\expandafter\@secondoftwo
\fi
{\expandafter\amsmath@bigm\csname fenced@\string#1\endcsname}%
{\amsmath@bigm#1}%
}
\newcommand{\DeclareFence}[2]{\@namedef{fenced@\string#1}{#2}}
\makeatother
\DeclareFence{\mid}{|}

\newcommand \vol{\mathrm{vol}}

\newcommand \id{\mathrm{id}}

\newcommand \F{\mathbb{F}}

\newcommand \R{\mathbb{R}}

\newcommand \Z{\mathbb{Z}}

\newcommand \SL{\mathrm{SL}}

\newcommand \GL{\mathrm{GL}}
\newcommand \PGL{\mathrm{PGL}}

\begin{document}

\title{Corrigendum and addendum to: Equidistribution of Eisenstein series in the level aspect}


\author{Ikuya Kaneko}
\address{Tsukuba Kaisei High School \\ 3315-10 Kashiwada, Ushiku, 300-1211 Japan}
\curraddr{}
\email{ikuyak@icloud.com}
\thanks{}

\author{Shin-ya Koyama}
\address{Department of Biomedical Engineering, Toyo University, 2100 Kujirai, Kawagoe, Saitama 350-8585 Japan}
\curraddr{}
\email{koyama@tmtv.ne.jp}
\thanks{}

\subjclass[2010]{Primary 58J51; Secondary 11F11, 11M12}

\keywords{}

\date{\today}

\dedicatory{}

\begin{abstract}
The second author formulated quantum unique ergodicity for Eisenstein series in the prime level aspect in ``Equidistribution of Eisenstein series in the level aspect'', Comm. Math. Phys. \textbf{289}, no. 3, 1131--1150 (2009). We point out major errors and propose ideas to correct particular parts of the proofs with partially weakened claims.
\end{abstract}

\maketitle

We shortly indulge in retrospection on variations of the \textit{quantum unique ergodicity} (QUE) conjecture put fourth by Rudnick and Sarnak~\cite{RudnickSarnak1994}. 
Luo and Sarnak~\cite{LuoSarnak1995} established the arithmetic QUE for the continuous spectrum spanned by Eisenstein series on $\SL_{2}(\Z) \backslash \mathbb{H}$. 
QUE in the level aspect was spelled out by Kowalski, Michel and VanderKam~\cite{KowalskiMichelVanderKam2002} for holomorphic cusp forms: they conjectured that the masses of newforms of fixed weight with large level $q$ are equidistributed amongst the fibers of the canonical projection $Y_{0}(q) \to Y_{0}(1)$ with $Y_{0}(q) = \Gamma_{0}(q) \backslash \mathbb{H}$ the modular curve. Nelson et al.~\cite{Nelson2011,NelsonPitaleSaha2014} have affirmatively answered their conjecture by following the strategy of Holowinsky and Soundararajan~\cite{HolowinskySoundararajan2010}. These types of variations intersect at second author's work~\cite{Koyama2009}, where QUE in the prime level aspect for Eisenstein series was focused.

In the introduction, the term of the quantum ergodicity has incorrectly used, and it refers to the equidistribution for a density one subsequence of eigenfunctions of the Laplacian within the unit cotangent bundle in the large eigenvalue limit. We wanted to generally mean QUE in configuration space. Moreover the coefficient in the right hand side of the first display on p.1132 should be $6/\pi$ in the light of Stade's formula on $\GL(2)$; for a correct statement, see~\cite[Eq. (7.9)]{HejhalRackner1992} or~\cite[Eq. (1.1)]{Spinu2003}. That is, one has
\begin{equation}\label{Luo-Sarnak}
\int_{\SL_{2}(\Z) \backslash \mathbb{H}} \phi(z) |E(z, 1/2+it)|^{2} \frac{dx dy}{y^{2}}
 = \frac{3}{\pi} \log(1/4+t^{2}) \int_{\SL_{2}(\Z) \backslash \mathbb{H}} \phi(z) \frac{dx dy}{y^{2}}+o(\log t)
\end{equation}
as $t \to \infty$ for a fixed smooth and compactly supported function $\phi: \SL_{2}(\Z) \backslash \mathbb{H} \to \R$ (cf.~\cite{Young2016} for a full main term with a power saving error term for~\eqref{Luo-Sarnak}). As a fatal matter, almost all the theorems and propositions should significantly be modified, since the claimed results are stronger than what is justified. In particular, we could no longer rely upon the Luo-Sarnak method when the level is varying. By gathering together various modifications described below, we see the correct version of Theorem 1.3 should morally be as follows:
\begin{theorem13}
Let $q$ be a prime or $q = 1$, and let $t$ to be a fixed real number. We denote by $\pi_{q}: Y_{0}(q) \to Y_{0}(1)$ the canonical projection with which fix a map $\iota_{q} = \iota$ such that $\pi_{q} \circ \iota = \id$ for each $q$. For any fixed compact Jordan measurable subsets $A, B \subset Y_{0}(1)$ having positive measure and for an arbitrary sequence in $\{(\iota_{q}, \kappa(q)) \mid \text{$\pi_{q} \circ \iota_{q} = \id$, $\kappa(q) = 0, i\infty$ for each $q$} \}$, we have
\begin{equation*}
\lim_{q \to \infty} 
\frac{\int_{A} |E_{q, \kappa(q)}(\iota z, 1/2+it)|^{2} d\mu}{\int_{B} |E_{q, \kappa(q)}(\iota z, 1/2+it)|^{2} d\mu}
 = \frac{\vol(A)}{\vol(B)},
\end{equation*}
with respect to the Poincar\'{e} measure $d\mu = y^{-2} dxdy$ on the modular curve $Y_{0}(1)$. 
\end{theorem13}
It should be mentioned that the assumption on $q$ could be weakened with more work, but at the very least we have to treat finitely many more cusps. The main issue in the original theorem was that the objects $A_{q}$ are too thin and the error terms that appeared in Propositions 1.4 and 1.5 are not provably smaller than the principal terms. It was also forgotten to ponder the error term emerging from the manipulation of the spectral decomposition of the characteristic function of a set $A_{q}^{0}$. Notice that $A_{q}^{0}$ is changing with $q$, so that such an error term depends on $q$. We are much indebted to Matthew Young and Jiakun Pan for their bringing those slips to our attention. As a proper generalization of~\cite{Koyama2009}, Jiakun Pan studied Eisenstein series with general nebentypus characters and general levels and succeeded in getting around the issues briefly described above. Apparently his strategy relies upon the generalization of Zagier's method of regularization to Eisenstein series on general levels; Zagier developed a formula for integrals of double and triple products of Eisenstein series on the level 1 case.

In order to circumvent many difficulties, it is conceivable to replace our chief concern with a softer object. Fortunately, the method we used works fine if instead of taking the thin sets of the form $A_{q}$, we select a set $A_{1}$ that is not dependent on $q$. Fix a nice test function $\phi$ on $\SL_{2}(\Z)$ and $\iota = \iota_{q}$ to denote a map which satisfies $\pi_{q} \circ \iota = \id$ for each $q$ (specifically, we have $\iota z \in Y_{0}(q)$ for $z \in Y_{0}(1)$). We thus consider the following ``unbalanced'' inner product against Eisenstein series on $\Gamma_{0}(q) \backslash \mathbb{H}$,
\begin{equation}\label{inner-product}
\langle |E|^{2}, \phi \rangle_{1} = \int_{Y_{0}(1)} \phi(z) |E_{q, \kappa(q)}(\iota z, 1/2+it)|^{2} \frac{dx dy}{y^{2}},
\end{equation}
although the work of Jiakun Pan is concerned with $\langle |E|^{2}, \phi \rangle_{q}$ for more general Eisenstein series like $E = E_{\infty}^{(q)}(z, s, \chi)$. We then spectrally decompose $\phi$ in~\eqref{inner-product} into cusp forms and incomplete Eisenstein series. The Luo-Sarnak approach proceeds by showing asymptotic formul{\ae} for the contributions of Maa{\ss} forms and incomplete Eisenstein series. Through working with~\eqref{inner-product}, there are no forms of level $q$ that enter after the application of the spectral decomposition, so that one could adapt the Luo-Sarnak method to address~\eqref{inner-product}. Actually, there is a technical difficulty in analyzing the integral handled in~\cite{Koyama2009} in the level aspect by our using only the spectral decomposition and Parseval's formula, since the Eisenstein series (absolute value squared) grows too fast at the cusp. As we gave heed before, it is adequate to think that renormalized integrals (such as in Michel and Venkatesh~\cite{MichelVenkatesh2010} using the language of representation theory) would be the most useful implement to sidestep the difficulty.

We also point out several problematic errors. However, the following corrections are meant for the original paper~\cite{Koyama2009} to have a mostly self-contained exposition, while they should be along with additional modifications when we consider~\eqref{inner-product}. First of all, throughout the paper the factor $n^{s-1}$ should be $|n|^{s-1}$ for $n \in \Z$. Furthermore, the condition on both sums in (2.3) and (2.4) should be recasted as $c \in \mathcal{C}(\kappa, \kappa^{\prime})$ for a pair of cusps $\kappa, \kappa^{\prime}$, where
\begin{equation*}
\mathcal{C}(\kappa, \kappa^{\prime}) = \bigg\{\gamma > 0: \begin{pmatrix} \ast & \ast \\ \gamma & \ast \end{pmatrix} \in 
\sigma_{\kappa}^{-1} \Gamma_{0}(q) \sigma_{\kappa^{\prime}} \bigg\}.
\end{equation*}
is the set of allowed moduli. By virtue of this, the change of variables $c = \gamma \sqrt{q}$ on p.1138 can clearly be justified. On p.1139, the definition of the Fourier coefficients $\rho_{j}(n)$ should be replaced with
\begin{equation*}
\rho_{j}(n) = 
	\begin{cases}
	q^{1/2} \tau_{j}(n/q) & \text{if $q \mid n$},\\
	0 & \text{otherwise}.
	\end{cases}
\end{equation*}
and this propagates typos throughout the later parts of the paper when we treat the oldforms of level 1 slashed by $q$. In particular, the assertion (2) in Lemma 2.4 is wrong and should be reformed to the shape that includes the extra $q^{1/2}$:
\begin{lemma24}[2]
When $u_{j}$ is an oldform for $\Gamma_{0}(q)$ and expressed by $u_{j}(z) = v_{j}(qz)$ with $v_{j}$ a cusp form for $\SL_{2}(\Z)$, we have
\begin{equation*}
\sum_{n = 1}^{\infty} \frac{\rho_{j}(n) \sigma_{\nu}(n)}{n^{s}}
 = \tau_{j}(1) q^{1/2-s} \frac{1+q^{\nu}-\tilde{\tau}_{j}(q) q^{\nu-s}}{1-q^{\nu-2s}} 
\frac{L(s, v_{j}) L(s-\nu, v_{j})}{\zeta(2s-\nu)}
\end{equation*}
and
\begin{equation*}
\sum_{n = 1}^{\infty} \frac{\rho_{j}(n) \sigma_{\nu}(nq^{-\alpha})}{n^{s}}
 = \tau_{j}(1) q^{1/2-s}(1+q^{\nu}) \frac{1-\tilde{\tau}_{j}(q) q^{\nu-s}}{1-q^{\nu-2s}} 
\frac{L(s, v_{j}) L(s-\nu, v_{j})}{\zeta(2s-\nu)}.
\end{equation*}
\end{lemma24}
Here we simplified the complicated numerator in (2.13) as well. The proof of this claim is always with the additional factor of $q^{1/2}$. In view of this correction, it is better to incorporate Lemma 2.6 with Lemma 2.5 to conclude the calculation of the contribution from cusp forms, because the proof of Lemma 2.5 includes incorrect statements (and that of Lemma 2.6 is back-of-the-envelope). For instance, on p.1143, the second-to-last line in the fourth display should be modified into
\begin{align*}
\frac{\rho_{j}(1) |L(s+it, u_{j})|^{2}}{\hat{\zeta}(1+2it) \zeta(2s)(1-q^{-2s})} 
\bigg(1-\frac{1-\tilde{\rho}_{j}(q) q^{-s-it}}{1-q^{-1-2it}} \bigg)\\
&\hspace{-6cm} = \frac{\rho_{j}(1) |L(s+it, u_{j})|^{2}}{\hat{\zeta}(1+2it) \zeta(2s)(1-q^{-2s})} 
\Big(1-(1-\tilde{\rho}_{j}(q) q^{-s-it})(1+q^{-1-2it}+\cdots) \Big)\\
&\hspace{-6cm} \ll \frac{\rho_{j}(q) q^{-s-it}}{1-q^{-2s}} 
\frac{|L(s+it, u_{j})|^{2}}{\hat{\zeta}(1+2it) \zeta(2s)} \ll q^{\frac{6\vartheta}{5}-\frac{1}{10}+\epsilon} \ll q^{\frac{1}{32}+\epsilon},
\end{align*}
where we exploited the bound in the recent work of Blomer, Humphries, Khan and Milinovich: $L(1/2, u_{j}) \ll q^{1/5+\vartheta/15+\epsilon}$ for $u_{j} \in \mathcal{B}^{\ast}(q)$ with $\mathcal{B}^{\ast}(q)$ being an orthonormal basis of Hecke-Maa{\ss} newforms of level $q$ having spectral parameter $t_{j}$. We denote by $\vartheta \in [0, 7/64]$ an admissible exponent towards the Ramanujan-Petersson conjecture and note that the appearance of $\vartheta$ also stems from the Fourier coefficients $\rho_{j}(q)$. Even if an optimal amplifier would yield the stronger bound $L(1/2, u_{j}) \ll_{\epsilon} q^{\frac{1}{4}-\frac{1-2\vartheta}{16}+\epsilon}$ of Burgess quality (that makes the limit of current technology), we cannot obtain any error term to be smaller than the main term in our QUE problem, unless we assume a slight improvement on the Ramanujan-Petersson conjecture. This error is irretrievable, nonetheless our dealing with the level aspect QUE problem in the shape~\eqref{inner-product} will be attended with no difficulties.

On the other hand, the full part of the last big display (2.15) on p.1143 is also mistaken. In order to rectify it, we use the recasted form of Lemma 2.4 described above, obtaining the following plausible estimate:
\begin{align*}
(2.15) &= \frac{1}{\hat{\zeta}(1+2it)} \bigg(\sum_{n = 1}^{\infty} \frac{\rho_{j}(n) \sigma_{-2it}(n)}{n^{s-it}}-\frac{1}{1-q^{-1-2it}} \sum_{n = 1}^{\infty} \frac{\rho_{j}(n) \sigma_{-2it}(nq^{-\alpha})}{n^{s-it}} \bigg)\\
& = \tau_{j}(1) \frac{q^{1/2+it-s}}{1-q^{-2s}} \frac{L(s+it, v_{j}) L(s-it, v_{j})}{\zeta(2s) \hat{\zeta}(1+2it)}\\
& \hspace{2cm} \times \Big(1+q^{-2it}-\tilde{\tau}_{j}(q) q^{-s-it}-(1+q^{-2it})
(1-\tilde{\tau}_{j}(q) q^{-s-it})(1+q^{-1-2it}+\cdots) \Big)\\
& = \tau_{j}(q) \frac{q^{1/2-2(s+it)}}{1-q^{-2s}} \frac{L(s+it, v_{j}) L(s-it, v_{j})}{\zeta(2s) \hat{\zeta}(1+2it)}
 + \text{(lower order terms)}\\
& \ll q^{\vartheta-1/2}.
\end{align*}
Hence, the contribution of the oldforms would not cause any problem in analyzing the integral $I_{j}(q, s)$. In any case our method of asymptotic evaluation for the inner product against incomplete Eisenstein series cannot work as long as we do not substitute $A_{1}$ for the thin sets $A_{q}$. So, it seems difficult to constructively approximate the projection of $\phi$ onto $\mathcal{E}(Y_{0}(q))$ in terms of incomplete Eisenstein series, where $\mathcal{E}(\Gamma \backslash \mathbb{H})$ is the orthogonal complement of the span of the Maass forms $\mathcal{C}(\Gamma \backslash \mathbb{H})$. In passing, the calculation of the contribution from incomplete Eisenstein series should also be significantly corrected, namely the factor 24 in the first term in Proposition 1.5 is incorrect. This error stems form a mis-calculation of a Dirichlet series counted with divisor functions, and a certain integral involving $K$-Bessel functions.

Although the work~\cite{Koyama2009} sticks to the $\GL(2)$ unitary Eisenstein series, it is also interesting to speculate what happens if we change the subject to higher ranks, such as to the $\GL(n)$ continuous spectrum spanned by the degenerate Eisenstein series induced from the maximal parabolic subgroup. Quite recently, Zhang~\cite{Zhang2019} studied an analogue of Luo and Sarnak's result in such direction, however QUE for higher rank non-degenerate Eisenstein series is much more difficult problem as it is closely related to the shifted convolution problem with generalized divisor functions and Fourier coefficients of higher rank Maass cusp forms. This way, it would also be possible to consider various versions of~\cite{Koyama2009} based solely on classical ideas of Luo and Sarnak. For example, the authors~\cite{KanekoKoyama2019} recently succeeded in showing equidistribution results for Eisenstein series in the level aspect \textit{over function fields}. In the case of the function field analogue, fundamental domains turn into Ramanujan graphs of the general form $\Gamma \backslash \PGL_{2}(F)/K$ with $F$ a local non-archimedean field, $K$ a maximal compact subgroup and $\Gamma$ a lattice in $\PGL_{2}(F)$. The setting we are interested in is the homogeneous space $G/K$, where $G = \PGL_{2}(k_{\infty})$ and $K = \PGL_{2}(r_{\infty})$ with $k_{\infty} = \F_{q}((1/t))$ the field of Laurent formal power series in the uniformizer $1/t$ over $\F_{q}$ and $r_{\infty}$ the ring of local integers $\F_{q}[[1/t]]$, respectively. Notice that $K$ is a maximal compact subgroup of $G$. Then the number of the discrete spectrum of the adjacency operator on $\mathcal{L}^{2}(\Gamma \backslash \mathcal{F})$ is known to be finite, where $\mathcal{F}$ is a $(q+1)$-regular tree. But from the viewpoint of the level aspect QUE problem, one can consider the Hecke congruence subgroup of the form $\Gamma_{0}(A)$ as $\deg A$ tends to infinity. The explicit forms of theorems in our work will be announced soon.

\providecommand{\bysame}{\leavevmode\hbox to3em{\hrulefill}\thinspace}
\providecommand{\MR}{\relax\ifhmode\unskip\space\fi MR }
\providecommand{\MRhref}[2]{%
  \href{http://www.ams.org/mathscinet-getitem?mr=#1}{#2}
}
\providecommand{\href}[2]{#2}

\end{document}